# Accretion vs. Core-Filament Collision

### v1

@@ 4.1 Accreting streamer or not?
Taniguchi et al. (2024) claimed that the streamer connects to the northern region through the bridge based on the integrated intensity map. They suggested that the streamer is accreting flow. However, as mentioned above, the gas in the bridge is likely to be different component from the streamer since they have different velocities: the bridge has the velocities from 6.5 (north) to 7.0 km s-1 (south) with the velocity gradient of 2.5 km s-1 pc-1 (hereafter, we define this direction is positive), whereas the streamer has the velocities from 7.5 (north) to 7.0 km s-1 (south) with the velocity gradient of –0.5 km s-1/0.05 pc = –10 km s-1 pc-1. There is no evidence showing that the streamer connects to the bridge in the velocity distribution.
In addition, the centre of gravity of the protostellar core has a significant offset from the protostar of 12″( au). If the streamer originates from the inflow from the ambient gas driven by the gravity, it is likely to point towards the core’s gravity centre (see Fig. 8). Or, if the offset is due to rotation, this object must have a large rotating envelope with a size of au. Such a structure is not detected towards the protostar (Tobin et al., 2016). This fact also implies that the streamer is not accretion origin, contrasting with the interpretation of Pineda et al. (2020). The northern cores (1&2) are close to gravitationally bound and therefore there is a possibility for them to evolve into protostars. Their merging might trigger star formation. The protostellar core also contains appreciable mass, a part of which may eventually accrete onto the central protostar before the ambient gas falls in.
Yano et al. (2024) proposed the DCC scenario to reproduce the density structure in Per-emb-2 in numerical simulations. According to this model, when two cores collide with each other, the shock compressed layer is formed in between, connected to the protostar. The gas is inflowing towards the protostar. However, if this is the case, both the bridge and streamer should have the velocity gradients with the same directions (see Kinoshita & Nakamura, 2022). The velocity/density structure appears more complicated than what the DCC model predicts. Therefore, the formation mechanism of the streamer is neither ambient gas inflow nor two round core collision.
@@ 4.2 A core-filament collision (CFC) model
From the evidence of the existence of the large scale filament, the positive gradient towards the streamer of the protostar. We propose a core-filament collision (CFC) model to explain for the configuration of Per-emb-2, Fig. 4 depicts the schematic representation of this CFC model.
@@
We propose that (1) a spherical core, with a relative speed of 1–1.5 km s-1, approached from the west side of the filament and collided with it. (2) The collision initiated star formation on the far side of the eastern extremity of the core, where the initial shock compression occurred, leading to the formation of a protostellar system. The arc-like streamer may be the bow structure formed by the CFC, and therefore the velocity gradient of the streamer is not due to the gravitational accretion. If its dynamics is not significantly affected by the stellar gravity, the flow is pointed towards outside along it (outflow), instead of infall. The bow typically has an almost linear velocity structure along it because of its curvature. Such a feature is consistent with the observation. When the stellar gravity is sufficiently strong to attract the gas towards the star, the flow may be pointing towards the protostar. It would be difficult to judge which direction (outflow or inflow) is reasonable from the current observational data. Future numerical experiments would be needed to verify this scenario more quantitatively.
@@
The head-tail morphology observed in the No. 3 core agrees well with the anticipated appearance resulting from this collision. The CCS arc (streamer) structure (Pineda et al., 2020) appears to well trace the edge of the protostellar core ( cm-2, see Fig. 2).
@@
Furthermore, the collision led to the division of the filament into two distinct structures (Nos. 1&2 and 5). The shock-compressed layers in the northern and southern regions attained densities of cm-3, resulting in relatively strong CCS/HC3N emissions. The northern shock-compressed layer became interconnected with the central protostellar system. There is a possibility that it could evolve into a accreting streamer in the future. However, the total accretion mass via the bridge (future streamer) may be small. The protostellar core envelope still contains enough material (). Therefore, the accretion along the streamers, if it exits, plays only a minor role in the mass accumulation towards the central protostar.
@@
The total velocity difference is about 1 (= 7.5–6.5) km s-1 and the filament width is about 0.1 pc, therefore it takes 0.1 pc / 1 km s-1 yr for the core to cross the filament. This is comparable to the typical lifetime of Class 0 protostar and consistent with the triggered star formation scenario proposed by CFC. Furthermore, if the bridge is the remnant of the shock compressed layer, its density increased from cm-3 to cm-3 by the isothermal shock with a Mach number of 1.5 km s-1 / sound speed . This condition in the shocked area, or streamer, is quite similar to the condition assumed by Suzuki et al. (1992) who started the chemical evolution calculation assuming a steady state with a uniform isothermal gas of cm-3 and K. According to Fig. 13 of (Suzuki et al., 1992), the CCS/HC3N becomes most abundant in yr in such a density range, and declines steeply after 0.5 Myr (see also Leung et al., 1984). These considerations on the timescales well match the CFC scenario. In our CFC or DCC scenarios, shocks are likely to somewhat increase gas temperature although the molecular gas is highly radiative and therefore the gas temperature is kept low even after shock compression. In such a situation, the streamers formed by the collision model may have emission of SO or CH3OH whose sublimation temperatures are about 50 and 80 K, respectively. These lines could exhibit broader widths as a result of shock interactions. The above consideration should be validated through future chemical evolution calculations using e.g., the Paris-Durham shock code (Godard et al., 2019).
@@ Appendix A Observations and target lines
The OTF observations were carried out in October 2022 (2 nights) and in March 2023 (1 daytime). Z45 is dual linear polarization receiver and we summed up the horizontal and vertical polarization components to improve the signal-to-noise ratios. The scan interval of the OTF observations was set to 8″, about a fifth of the beam size. The pointing observations were made every one hour with SiO maser lines of NML-Tau (KL Tau). The pointing errors were within 3″– 5″. We used SAM45 digital spectrometer as a backend. SAM45 is a highly flexible FX-type digital spectrometer with a finest frequency resolution of 3.81 kHz. The bandwidth of SAM45 with 3.81 kHz was 15.625 MHz. It took about 30 min to obtain a single map. We summed 6 maps to obtain a final map. The total observation time was about 6 hours including the overhead of pointing observations. The on-source time was 3 hours. During the observations, the typical system noise temperatures were about 150–200 K.
The rest frequencies and transitions of the target lines are summarised in Table 2. See Nakamura et al. (2024) for the rest frequency of CCS.
@@ B.1 Core identification
We applied astrodendro (Rosolowsky et al. 2008) to the high-resolution Herschel column density map with an angular resolution of 18″. The astrodendro has three parameters for the hierarchical structure analysis: min_value which is the threshold value, min_delta which is the minimum step for structure identification, and min_npix which is the minimum pixel number. We adopt min_value , min_delta , and min_npix = Beam size/pixel size, cm-2) is the rms noise level. The astrodendro identifies three structures: a leaf, branch, and trunk. The minimum structure of “a leaf” is defined as a dense core. To estimate the ambient gas pressure around the cores, we adopted the physical quantities averaged in a branch immediately below the corresponding leaf. We also derive the core mass by subtracting the background component which is defined as the trunk.
@@ B.2 Dynamical states of the cores
and is the gravitational constant, is the Boltzman constant, is the velocity dispersion measured by CCS, (=2.33) is the mean molecular weight, is the mass of a hydrogen atom, and is the mass of a CCS molecule. The CCS velocity dispersions towards the cores are computed by the Gaussian fitting of the averaged spectra (see Fig. 5). It is worth noting that CCS emission may preferentially trace the outer envelopes of cores, given its critical density of cm-3. In contrast, the core densities calculated from the Herschel image range from to cm-3 as shown in Table 1. Despite this, in the core analysis presented here, we adopt the CCS line widths as representative of the cores’ properties.
@@
where and are the density and 1D intrinsic velocity dispersion of ambient gas, respectively, and they are derived from the values averaged in the branch immediately below the corresponding leaf. The ambient density was computed under the assumption that the branch is spherical symmetry.
The core virial parameter is calculated as
Table 3 summarises the velocity dispersions of the cores and their energy ratios. Fig. 6 shows the energy ratios vs. , where W, U, and S are the gravity term, internal kinetic and thermal energy term, and the surface term, respectively. In this plot, the cores below the black dashed line are satisfied with the condition , for which the cores collapse. The core’s self-gravity is important for the area above the dotted line (), whereas the surface pressure is dominant in the area below the dotted line. The cores 1, 2, and 3 are located in the gravity-dominant area. They are close to the dynamical equilibrium and self-gravitating. For the core 5, the external pressure plays a crucial role in the core’s dynamical stability, and thus the core is dynamically unstable due to the external pressure. Such a core is difficult to create from the gravitational fragmentation of the filament. The formation of unstable non-self-gravitating core (core No. 5) appears to agree with the collision scenario.
@@ Appendix C Channel maps
Fig. 7 shows the channel maps of HC3N ( and HC5N (). The HC3N has a stronger emission but having several overlapped hyperfine components. However, adopting the rest frequency of as a centre frequency, we simply integrated all the emission to construct the channel maps. The bridge is prominent for HC3N (, and structure similar to CCS was detected. For HC5N (, the strong emission is seen just at the southern edge of cores No. 1 and 2.
@@
1 | 0.37 | 0.41 | 0.26 | 0.28 | -0.703 | -0.111 | 1.422 | |
2 | 0.41 | 0.41 | 0.27 | 0.27 | -0.986 | -0.019 | 1.014 | |
3 | 0.84 | 0.61 | 0.41 | 0.33 | -0.928 | -0.734 | 1.077 | |
5 | 0.45 | 0.61 | 0.28 | 0.34 | -0.339 | -3.025 | 2.954 |
Appendix D Centre of gravity of the protostellar core w.r.t the large filament
@@
If the curved shape of the streamer identified by Pineda et al. (2020) is due to the angular momentum of infalling gas, the rotational motion must be significant on the scale of au, leading to the formation of a large ( au) rotating envelope or disk. However, such a large circumstellar structure was not detected in the observations by Tobin et al. (2016). Thus, at this scale, the rotational motion would still be minor and unlikely to create the elongated, curved structure seen in Fig. 2. For example, Yano et al. (2024)’s core collision simulation showed that the streamer formed by the infalling gas from the second core is smoothly connected to a central, rotating disk-like envelope. Such a structure is a plausible outcome of the infalling gas scenario. Since the rotating disk is small for this protostar, the curved structure, if formed by the infalling gas, must almost point directly to the protostar.

### v2

@@ 4.1 Accreting streamer or not?
Taniguchi et al. (2024) claim, based on the integrated intensity map, that the streamer connects to the northern region through the bridge. They suggest that the streamer is an accreting flow. However, as mentioned above, the gas in the bridge is likely a separate component since the streamer and bridge have different velocities: the velocity of the bridge ranges from 6.5 (in the north) to 7.0 km s-1 (south) with a velocity gradient of 2.5 km s-1 pc-1 (hereafter, we define this direction as positive), whereas the streamer has velocities of 7.5 (north) to 7.0 km s-1 (south) with a velocity gradient of –0.5 km s-1/0.05 pc = –10 km s-1 pc-1. From on the velocity distribution, there is no evidence that the streamer connects to the bridge.
In addition, the centre of gravity of the protostellar core is significantly offset from the protostar, by 12″( au). If the streamer originates from the inflow from the ambient gas driven by the gravity, it likely points towards the core’s gravity centre (see Fig. 8). If the offset were due to rotation, this object would have a large rotating envelope with a size of au. Such a structure is not detected towards the protostar (Tobin et al., 2016). This also implies that the streamer is not an origin of accretion, contrasting with the interpretation of Pineda et al. (2020). The northern cores (1 and 2) are close to gravitationally bound and therefore could potentially evolve into protostars. Their merging could trigger star formation. The mass of the protostellar core is significant; a part of it may eventually accrete onto the central protostar before the ambient gas falls in.
Yano et al. (2024) used the DCC scenario to reproduce the density structure in Per-emb-2 in numerical simulations. According to this model, when two cores collide with each other, a shock-compressed layer is formed in between them and connects to the protostar. The gas is inflowing towards the protostar. However, if this is the case, both the bridge and streamer should have velocity gradients in the same direction (see Kinoshita & Nakamura, 2022). The velocity and density structures appear more complicated than what the DCC model predicts. Therefore, the formation mechanism of the streamer is neither ambient gas inflow nor two-round core collision.
@@ 4.2 A core-filament collision (CFC) model
Based on the evidence of the existence of the large-scale filament and a positive gradient towards the streamer of the protostar, we propose a CFC model to explain the configuration of Per-emb-2. Figure 4 depicts the schematic representation of this CFC model.
@@
We propose that: (1) A spherical core, with a relative speed of 1–1.5 km s-1, approached from the west side of the filament and collided with it. (2) The collision initiated star formation on the far side of the eastern extremity of the core, where the initial shock compression occurred, leading to the formation of a protostellar system. The arc-like streamer may be the bow structure formed by the CFC, and therefore the velocity gradient of the streamer is not due to gravitational accretion. If its dynamics are not significantly affected by the stellar gravity, the flow along the streamer should be pointed towards the outside (i.e. it is an outflow and not infalling). The bow typically has an almost linear velocity structure because of its curvature. Such a feature is consistent with the observation. When the stellar gravity is sufficiently strong to attract the gas towards the star, the flow may be pointing towards the protostar. Since it is difficult to judge the direction (outflow or inflow) from the current observational data, numerical experiments are needed to verify this scenario more quantitatively.
@@
The head-tail morphology observed in core 3 agrees well with what would result from such a collision. The CCS arc (streamer) structure (Pineda et al., 2020) appears to trace the edge of the protostellar core ( cm-2; see Fig. 2).
@@
Furthermore, the collision led to the division of the filament into two distinct structures (cores 1 and 2 and core 5). The shock-compressed layers in the northern and southern regions attained densities of cm-3, resulting in relatively strong CCS and HC3N emissions. The northern shock-compressed layer became interconnected with the central protostellar system. There is a possibility that it could evolve into an accreting streamer in the future. However, the total accretion mass via the bridge (future streamer) may be low. The protostellar core envelope still contains enough material (). Therefore, the accretion along the streamers, if it exits, plays only a minor role in the mass accumulation towards the central protostar.
@@
The total velocity difference is about 1 (= 7.5–6.5) km s-1, and the filament width is about 0.1 pc; therefore, it takes 0.1 pc / 1 km s-1 yr for the core to cross the filament. This is comparable to the typical lifetime of a Class 0 protostar and is consistent with the triggered star formation proposed in the CFC model scenario. Furthermore, if the bridge is the remnant of the shock-compressed layer, its density must have increased from cm-3 to cm-3 due to the isothermal shock with a Mach number of 1.5 km s-1/sound speed . This condition in the shocked area, or streamer, is similar to the condition assumed by Suzuki et al. (1992), who started the chemical evolution calculation assuming a steady state with a uniform isothermal gas of cm-3 and K. According to Fig. 13 of Suzuki et al. (1992), CCS and HC3N become most abundant in yr in such a density range, and declines steeply after 0.5 Myr (see also Leung et al., 1984). These timescales match the CFC scenario. In our CFC or DCC scenarios, shocks are likely to increase the gas temperature somewhat, but the molecular gas is highly radiative and therefore the gas temperature stays low even after shock compression. In such a situation, the streamers formed by the collision model may have emission of SO or CH3OH, whose sublimation temperatures are about 50 and 80 K, respectively. These lines could exhibit broader widths as a result of shock interactions. This hypothesis should be validated with chemical evolution calculations using, for example, the Paris-Durham shock code (Godard et al., 2019).
@@ Appendix A Observations and target lines
The on-the-fly observations were carried out in October 2022 (2 nights) and in March 2023 (1 daytime). Z45 is dual linear polarisation receiver and we summed up the horizontal and vertical polarisation components to improve the signal-to-noise ratios. The scan interval of the on-the-fly observations was set to 8″, about a fifth of the beam size. The pointing observations were made every one hour with SiO maser lines of NML-Tau (KL Tau). The pointing errors were within 3″– 5″. We used SAM45 digital spectrometer as a backend. SAM45 is a highly flexible FX-type digital spectrometer with a finest frequency resolution of 3.81 kHz. The bandwidth of SAM45 with 3.81 kHz was 15.625 MHz. It took about 30 min to obtain a single map. We summed 6 maps to obtain a final map. The total observation time was about 6 hours including the overhead of pointing observations. The on-source time was 3 hours. During the observations, the typical system noise temperatures were about 150–200 K.
The rest frequencies and transitions of the target lines are summarised in Table 2 (see Nakamura et al. 2024 for the rest frequency of CCS).
@@ B.1 Core identification
We applied astrodendro (Rosolowsky et al. 2008) to the high-resolution Herschel column density map with an angular resolution of 18″. The astrodendro has three parameters for the hierarchical structure analysis: min_value which is the threshold value, min_delta which is the minimum step for structure identification, and min_npix which is the minimum pixel number. We adopt min_value , min_delta , and min_npix = beam area/pixel size, cm-2) is the rms noise level. The astrodendro identifies three structures: a leaf, branch, and trunk. The minimum structure of ‘a leaf’ is defined as a dense core. To estimate the ambient gas pressure around the cores, we adopted the physical quantities averaged in a branch immediately below the corresponding leaf. We also derive the core mass by subtracting the background component which is defined as the trunk.
@@ B.2 Dynamical states of the cores
and is the gravitational constant, is the Boltzmann constant, is the velocity dispersion measured by CCS, (=2.33) is the mean molecular weight, is the mass of a hydrogen atom, and is the mass of a CCS molecule. The CCS velocity dispersions towards the cores are computed by the Gaussian fitting of the averaged spectra (see Fig. 5). It is worth noting that CCS emission may preferentially trace the outer envelopes of cores, given its critical density of cm-3. In contrast, the core densities calculated from the Herschel image range from to cm-3 as shown in Table 1. Despite this, in the core analysis presented here, we adopt the CCS line widths as representative of the cores’ properties.
@@
where and are the density and 1D intrinsic velocity dispersion of ambient gas, respectively, and they are derived from the values averaged in the branch immediately below the corresponding leaf. The ambient density was computed under the assumption that the branch is spherical symmetry. The core virial parameter is calculated as
Table 3 summarises the velocity dispersions of the cores and their energy ratios. Figure 6 shows the energy ratios versus , where W, U, and S are the gravity term, internal kinetic and thermal energy term, and the surface term, respectively. In this plot, the cores below the black dashed line are satisfied with the condition , for which the cores collapse. The core’s self-gravity is important for the area above the dotted line (), whereas the surface pressure is dominant in the area below the dotted line. The cores 1, 2, and 3 are located in the gravity-dominant area. They are close to the dynamical equilibrium and self-gravitating. For the core 5, the external pressure plays a crucial role in the core’s dynamical stability, and thus the core is dynamically unstable due to the external pressure. Such a core is difficult to create from the gravitational fragmentation of the filament. The formation of unstable non-self-gravitating core (core 5) appears to agree with the collision scenario.
@@ Appendix C Channel maps
Figure 7 shows the channel maps of HC3N ( and HC5N (). The HC3N has a stronger emission but having several overlapped hyperfine components. However, adopting the rest frequency of as a centre frequency, we simply integrated all the emission to construct the channel maps. The bridge is prominent for HC3N (, and structure similar to CCS was detected. For HC5N (, the strong emission is seen just at the southern edge of cores 1 and 2.
@@
| 1 | 0.37 | 0.41 | 0.26 | 0.28 | -0.703 | -0.111 | 1.422 | |
| 2 | 0.41 | 0.41 | 0.27 | 0.27 | -0.986 | -0.019 | 1.014 | |
| 3 | 0.84 | 0.61 | 0.41 | 0.33 | -0.928 | -0.734 | 1.077 | |
| 5 | 0.45 | 0.61 | 0.28 | 0.34 | -0.339 | -3.025 | 2.954 |
Appendix D Centre of gravity of the protostellar core with respect to the large filament
@@
If the curved shape of the streamer identified by Pineda et al. (2020) is due to the angular momentum of infalling gas, the rotational motion must be significant on the scale of au, leading to the formation of a large ( au) rotating envelope or disk. However, such a large circumstellar structure was not detected in the observations by Tobin et al. (2016). Thus, at this scale, the rotational motion would still be minor and unlikely to create the elongated, curved structure seen in Fig. 2. For example, Yano et al. (2024)’s core collision simulation showed that the streamer formed by the infalling gas from the second core is smoothly connected to a central, rotating disk-like envelope. Such a structure is a plausible outcome of the infalling gas scenario. Since the rotating disk is small for this protostar, the curved structure, if formed by the infalling gas, must almost point directly to the protostar. In Fig. 8 we have also indicated the position of the filament shown in Fig. 4.